ELECTRONIC SPECTRA OF METAL-DILIGAND COMPLEXES: Cu(NH_{3})_{2} AND Cu(ND_{3})_{2}

\underline{SHENGGANG LI}, BRADFORD R. SOHNLEIN, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Zero electron kinetic energy (ZEKE) photoelectron spectroscopy was used to study the electronic spectra of metaldiammonia and its deuterated species. The adiabatic ionization potential of the copper complex was measured to be 29532 cm^{−1}, a shift of more than a factor of two relative to that of the copper atom (62317 cm^{−1}); deuteration further decreased the ionization potential of the complex by 219 cm^{−1}due to the shift of the zero point vibration energies. Furthermore, the ZEKE spectra revealed the copper-ammonia vibrations and ammonia torsion. For example, in Cu(NH_{3})_{2} the metal ion -ligand symmetric stretching and bending frequencies were measured to be 436 and 139 cm^{−1}, whereas the NH_{3} torsion was determined as 32 cm^{−1}.